John has three sacks of apples and three more apples in his pocket. Each sack contains the same number of apples. Altogether, John has 33 apples. How many apples are in each sack?
Answer by checkley71(8403)

3 EQUAL SACKS = 3X
PLUS THE 3 IN HIS POCKET IS
3X+3=33
3X=33-3
3X=30
X=30/3
X=10 APPLES IN EACH OF THE 3 SACKS.